ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

4. Algebraic Geometry

1991 MS Classification: 14E40 32S15 Kwieciński, Michał , Uniwersytet Jagielloński, Kraków (Poland) An algebraic map with nonconstructible topology We construct a simple example of a complex algebraic map, which (even locally) has an uncountable number of local topological types at points of the source space. This contrasts with the local topology of complex algebraic spaces where, by the theory of Whitney stratifications, one has a locally finite number of local topological types. Reference: M. Kwieciński, A complex map with complex topology, Proc. of the AMS, to appear